# Coarse-to-Fine Grounded Memory for LLM Agent Planning

### v1

@@ 4.2 Main Results
Effectiveness of CFGM. As shown in Table 1, CFGM outperforms ReAct across all environments: AlfWorld SR (+10.40%), WebShop SR (+20%), and ScienceWorld SR (+31%). This performance boost is due to CFGM’s coarse-to-fine grounding approach, which actively integrates LLM knowledge into structured memories. Compared to other memory-augmented agent baselines, our method also achieves superior task performance. This improvement stems from two key aspects: firstly, our approach leverages both offline experiences from long-term memory to assist online planning and incorporates short-term memory reflection on current trajectories; secondly, we utilize the internal knowledge of LLMs to ground the memories, enabling the agent to better comprehend and utilize past experiences.
@@ 4.3 Ablation Studies
After evaluating the overall performance of the framework, we conduct extensive experiments to assess the effectiveness of our framework and the design of memory grounding process of each component.
@@
Impact of Experience Retrieval Scope on Tips Quality. For the hybrid-grained tips used online, we investigate how the number of retrieved experiences (top-$k$) affects the tips’ quality. Specifically, we evaluate the impact by the SR on AlfWorld using a ReAct + ET setup. As shown in 3, the baseline ReAct ($k=0$) has an SR of 80.60%. Introducing just one relevant experience’s tips ($k=1$) improves SR to 82.09%. Performance peaks at $k=2$ and $k=3$, both achieving 86.57% SR. Increasing k further to 5 leads to a drop in performance (81.34%), due to the inclusion of less relevant or potentially conflicting hybrid-grained tips. This suggests that retrieving a moderate, focused set of the most relevant experiences ($k=2$ or $3$) provides the highest quality hybrid-grained tips to guide the LLM agent.
@@
Comparative Analysis of Online Reflection. To explore the rationality of KIR’s memory grounding design, we analyze the effectiveness of different online reflection mechanisms on AlfWorld. Table 5 compares KIR against two other reflection variants: a template-free Self-QA reflection without key information and a QA reflection via fixed questioning template like QuBE (Kim et al., 2024)), both applied on top of a ReAct agent. As shown in Table 5, KIR demonstrates superior corrective assistance for agent planning compared to QA Reflection (Fixed Question), which indicates that fine-grained grounded memory can achieve more effective reflection outcomes with greater flexibility than manually human-crafted question-answering reflection approaches. The performance improvement of KIR over Self-QA Reflection further demonstrates that prior analysis of short-term memory can significantly enhance subsequent comprehension and reasoning effectiveness when utilizing grounded memory.
@@ Appendix A Environment Details
A.4 WebArena-Shopping (Zhou et al., 2023)
WebArena is a suite of realistic and reproducible web environments designed for the development and evaluation of autonomous agents. The platform features fully operational, self-hostable web applications that emulate real-world websites across various domains.The WebArena–Shopping environment is implemented as part of the WebArena suite, providing a realistic and reproducible e-commerce setting for interactive planning agents.
@@ Appendix B More Implementation Details
We adopted different experimental task selection strategies for different environments. In the case of ALFWorld (Shridhar et al., 2020), we utilized the 134 solvable tasks that ReAct and ExpeL used, which comprised of 6 different task types in a virtual household environment: Pick (24), Clean (18), Heat (31), Cool (23), Examine (21), and Pick two (17). Similarly, for WebShop Yao et al. (2022), we evaluated using the same 100 tasks used by ReAct and ExpeL. For ScienceWorld (Wang et al., 2022), we assess performance using 100 tasks sampled from test set, which comprised of diverse task types with environmental steps up to 80: task-10-use-thermometer (10), task-2-power-component (5), task-2a-test-conductivity(10), task-3-find-animal (10), task-3-find-living-thing (10), task-3-find-non-living-thing (10), task-3-find-plant (10), task-4-grow (8), task-5-chemistry (11), task-6-lifespan (3), task-7-identify-life-stage (9), task-1-change-the-state-of-matter (4).

### v2

@@ 4.2 Main Results
Effectiveness of CFGM. As shown in Table 1, CFGM outperforms ReAct across all environments: AlfWorld SR (+10.40%), WebShop SR (+20%), and ScienceWorld SR (+31%). This performance boost is due to CFGM’s coarse-to-fine grounding approach, which actively integrates LLM knowledge into structured memories (zhang2023replay). Compared to other memory-augmented agent baselines, our method also achieves superior task performance. This improvement stems from two key aspects: firstly, our approach leverages both offline experiences from long-term memory to assist online planning and incorporates short-term memory reflection on current trajectories; secondly, we utilize the internal knowledge of LLMs to ground the memories, enabling the agent to better comprehend and utilize past experiences. Additionally, CFGM consistently and significantly outperforms the ExpeL + QuBE baseline. This demonstrates CFGM’s superiority is not merely an artifact of combining two orthogonal techniques.Furthermore, we wish to re-emphasize that CFGM is not a simple combination of existing ideas. Its components are deeply integrated and synergistic with coarse-to-fine grounded memory. Specifically, for example, the coarse-grained Focus Points are not just an independent pre-processing step; they fundamentally improve the quality of the memories collected. This, in turn, enhances the effectiveness of both the hybrid-grained Experience-wise Tips and the fine-grained Key Information Reflection. This synergistic "1+1>2" effect is a core contribution of our work and distinguishes CFGM from a simple integrated pipeline of unrelated modules.
@@ 4.3 Ablation Studies
After evaluating the performance of the framework, we conduct extensive experiments to assess the framework’s effectiveness and the design of memory grounding process of each component.
@@
Impact of Experience Retrieval Scope on Tips Quality. For the hybrid-grained tips used online, we investigate how the number of retrieved experiences (top-$k$) affects the tips’ quality. Specifically, we evaluate the impact by the SR on AlfWorld using a ReAct + ET setup. As shown in Table 3, the baseline ReAct ($k=0$) has an SR of 80.60%. Introducing just one relevant experience’s tips ($k=1$) improves SR to 82.09%. Performance peaks at $k=2$ and $k=3$, both achieving 86.57% SR. Increasing k further to 5 leads to a drop in performance (81.34%), due to the inclusion of less relevant or potentially conflicting hybrid-grained tips. This suggests that retrieving a moderate, focused set of the most relevant experiences ($k=2$ or $3$) provides the highest quality hybrid-grained tips to guide the LLM agent.
@@
Comparative Analysis of Online Reflection. To explore the rationality of KIR’s memory grounding design, we analyze the effectiveness of different online reflection mechanisms on AlfWorld. Table 5 compares KIR against two other reflection variants: a template‑free Self-QA reflection without key information and a QA reflection via fixed questioning template like QuBE (Kim et al., 2024)), both applied on top of a ReAct agent. As shown in Table 5, KIR demonstrates superior corrective assistance for agent planning compared to QA Reflection (Fixed Question), which indicates that fine-grained grounded memory can achieve more effective reflection outcomes with greater flexibility than manually human-crafted question-answering reflection approaches. The performance improvement of KIR over Self-QA Reflection further demonstrates that prior analysis of short-term memory can significantly enhance subsequent comprehension and reasoning effectiveness when utilizing grounded memory.
@@
The Overall Efficiency of CFGM. To assess the practical efficiency of CFGM compared against the baseline methods, we analyzed the token consumption and interaction turns of the framework on the AlfWorld benchmark. As shown in Table 6, during offline processing, CFGM’s token cost is comparable to ExpeL and notably more efficient than AutoGuide, which show the efficient offline traininng process of CFGM. Additionally, in the online phase, CFGM is the most efficient framework among the memory-augmented baselines in terms of both tokens per step and average interaction turns. This reduction in steps can lead to lower overall latency and cost, demonstrating that the comprehensive memory grounding leads to more efficient, less redundant agent behaviors.
Method Off. Tokens On. Tokens On. Turns
ReAct - 1734.6 19.01
ExpeL 3888.2 5125.5 17.32
AutoGuide 4873.4 4809.4 16.55
QuBE - 4752.3 17.21
CFGM 4068.5 4681.4 14.32
Table 6: Average token consumption for processing a single training data sample during the offline phase on AlfWorld (Off. Tokens). Average token consumption per step and average interaction turns per complete trajectory during online inference on AlfWorld (On. Tokens and On. Turns).
@@ 5 Conclusion
Acknowledgments
This work is supported by the National Key R&D Program of China (No.2022ZD0116405).
@@ A.2 WebShop Yao2022
A.4 WebArena-Shopping (zhou2023webarena)
WebArena is a suite of realistic and reproducible web environments designed for the development and evaluation of autonomous agents. The platform features fully operational, self-hostable web applications that emulate real-world websites across various domains.The WebArena–Shopping environment is implemented as part of the WebArena suite, providing a realistic and reproducible e‑commerce setting for interactive planning agents.
@@ Appendix B More Implementation Details
We adopted different experimental task selection strategies for different environments. In the case of ALFWorld (Shridhar2020), we utilized the 134 solvable tasks that ReAct and ExpeL used, which comprised of 6 different task types in a virtual household environment: Pick (24), Clean (18), Heat (31), Cool (23), Examine (21), and Pick two (17). Similarly, for WebShop Yao2022, we evaluated using the same 100 tasks used by ReAct and ExpeL. For ScienceWorld (Wang et al., 2022), we assess performance using 100 tasks sampled from test set, which comprised of diverse task types with environmental steps up to 80: task-10-use-thermometer (10), task-2-power-component (5), task-2a-test-conductivity(10), task-3-find-animal (10), task-3-find-living-thing (10), task-3-find-non-living-thing (10), task-3-find-plant (10), task-4-grow (8), task-5-chemistry (11), task-6-lifespan (3), task-7-identify-life-stage (9), task-1-change-the-state-of-matter (4).